OpenStudy (anonymous):

davante wrote the equation y=25-5x to describe his distance from home, in miles, after x minnutes of walking. how far from home was Danvante when he began walking home?

OpenStudy (sethzk):

Im here to help

OpenStudy (anonymous):

hiiiii

OpenStudy (sethzk):

Do you still need help?

OpenStudy (sethzk):

??

OpenStudy (anonymous):

yes please :D

OpenStudy (sethzk):

ok Lets identify the varibles. Can you tell me what y represents?

OpenStudy (anonymous):

i think the y represent how far from home when D began walk home. :D

OpenStudy (sethzk):

correct

OpenStudy (sethzk):

Now tell me what x represents

OpenStudy (sethzk):

nvm

OpenStudy (anonymous):

ok, though just to be sure i think the x represents the minutes of walking :D

OpenStudy (sethzk):

5 is the rate at which he is walking

OpenStudy (anonymous):

ok....

OpenStudy (sethzk):

x is how long he has been walking. so to find how far he has walked you multiply the two (5 and x). Now to find the distance from home you subtract the product of 5 and x from y

OpenStudy (anonymous):

ok, so 5x, x is the minute after he walked....so 5x is the total minute he walked right?

OpenStudy (sethzk):

no. 5x is how far he has walked

OpenStudy (sethzk):

is the question multiple choice?

OpenStudy (anonymous):

no, sir ( also sorry cuz i didn't see ur post becuz my computer is lagging) :D

OpenStudy (sethzk):

It is ok

OpenStudy (sethzk):

Well can I just tell you the answer?

OpenStudy (anonymous):

yes please but can i try to answer it to mak sure i understand

OpenStudy (sethzk):

Yes

OpenStudy (sethzk):

By all means

OpenStudy (mathmale):

Suppose that this guy has not yet begun walking. How far is he from home at that time?

OpenStudy (sethzk):

OH I know how to explain it in a much better way now

OpenStudy (sethzk):

MathMale wanna check out my question?

OpenStudy (anonymous):

Can you show me how to perform the problem to find the distance. so far i got y-(5*x-25)=0 but i'm not sure :(

OpenStudy (sethzk):

If devante has been walking for 0 minutes write the original equation like that

OpenStudy (sethzk):

where x=0

OpenStudy (sethzk):

the distance from home is 25

OpenStudy (sethzk):

because y=25-5x if x=0 tjen it becomes y=25

OpenStudy (sethzk):

and y=distance from home

OpenStudy (sethzk):

Medal?

OpenStudy (anonymous):

ohhh, THank you so much, sir.

OpenStudy (sethzk):

Noooooooo problem